OpenStudy (anonymous):

A number that multiplies to -12 but adds to -5

OpenStudy (anonymous):

that is a hell of a number!

OpenStudy (anonymous):

numbers that multiply to -12: 4 * -3 = -12 -4 * 3 = -12 1 * -12 = -12 -1 * 12 = -12 2 * -6 = -12 -2 * 6 = -12

OpenStudy (anonymous):

but i suppose it is really "two numbers whose product is -12 and whose sum is -5 you don't have that many whole number choices

OpenStudy (anonymous):

yeah what i replied was whole numbers :)

OpenStudy (anonymous):

solve \[x^2-5x-12=0\]

OpenStudy (anonymous):

yep ^ bingo :)

OpenStudy (anonymous):

OK thanks :)